OpenStudy (anonymous):

how do you solve 9*(-2 1/10)

OpenStudy (anonymous):

could you show me step by step on how you got that answer please

OpenStudy (angela210793):

w8 a sec pls..i may have done smth wrong

OpenStudy (angela210793):

is it -2*1/10 or -21/10?

OpenStudy (anonymous):

9* -2 1/10 nine times negative 2 and one tenth

jhonyy9 (jhonyy9):

-18/10 = -9/5

OpenStudy (angela210793):

-2*1/10=(-2*10+1)/10=-19*10 sp 9*-19/10=-171/10

OpenStudy (anonymous):

so the answer is .............

OpenStudy (angela210793):

tht's wht i got....idk why Jhonny got another 1 O.o

OpenStudy (anonymous):

what was your answer anglea

OpenStudy (amistre64):

just multiply

OpenStudy (angela210793):

9*-19/10=-171/10 tht's my answer.....

OpenStudy (amistre64):

-18' 1/10

OpenStudy (amistre64):

can we do that?

OpenStudy (amistre64):

9(-2 -1/10) -18' -9/10

OpenStudy (amistre64):

-18 -9/10 = N ; *10 -180 -9 = 10N ; add to add -189 = 10N ; /10 -189/10 = N

OpenStudy (anonymous):

angela you're mixing up products and sums it should be like this: 9* -2 1/10 nine times negative 2 and one tenth in case the formula is: 9 * (-2) * 1/10 9 * (-2) * 1/10 = -18 * 1/10 = -18/10 = -9/5 = -1,8 in case the formula is: 9 * (-2 + 1/10) 9 * (-2 + 1/10) = 9* (-20/10 + 1/19) = 9* (-19/10) = -171/10 = -17,1

OpenStudy (amistre64):

-2' 1/10 is actually the sum of: (-2 - 1/10)